# Zechen Li1,  Shohreh Deldari1,  Linyao Chen2,  Hao Xue1,  Flora D. Salim2

### v2

@@ Abstract
We introduce SensorLLM, a two-stage framework that enables Large Language Models (LLMs) to perform human activity recognition (HAR) from sensor data. Despite their strong reasoning and generalization capabilities, LLMs remain underutilized for motion sensor data due to the lack of semantic context in time-series, computational constraints, and challenges in processing numerical inputs. SensorLLM addresses these limitations through a Sensor-Language Alignment stage, where we introduce special tokens for each sensor channel and automatically generate textual trend descriptions. This alignment enables LLMs to capture numerical variations, channel-specific features, and data of varying durations—without requiring human annotations. In the subsequent Task-Aware Tuning stage, we refine the model for HAR classification, achieving performance that matches or surpasses state-of-the-art methods. Our results demonstrate that SensorLLM evolves into an effective sensor learner, reasoner, and classifier through Sensor-Language Alignment, generalizing across diverse HAR datasets. We believe this work establishes a foundation for future research on time-series and text alignment, paving the way for foundation models in sensor data analysis. Our codes are available at https://github.com/zechenli03/SensorLLM.
[Uncaptioned image] SensorLLM: Aligning Large Language Models with Motion Sensors for Human Activity Recognition
@@ 1 Introduction
Human Activity Recognition (HAR) is a time-series classification task that maps sensor signals, such as accelerometer and gyroscope data, to human activities. Traditional models like LSTM (Guan and Plötz, 2017; Hammerla et al., 2016) and DeepConvLSTM (Ordóñez and Roggen, 2016) learn high-level features but are task-specific and struggle to generalize across different sensor configurations and activity sets. In contrast, Large Language Models (LLMs) (Han et al., 2021) have shown remarkable success in integrating diverse data types (Wu et al., 2023b; Yin et al., 2023), including text and images.
@@
Enabling LLMs to process sensor data (Jin et al., 2023) requires either (1) pretraining or fine-tuning on time-series data (Zhou et al., 2023a), which demands substantial computational resources and is hindered by limited and imbalanced labeled data, or (2) leveraging zero-shot and few-shot prompting by converting sensor data into text (Kim et al., 2024; Ji et al., 2024). The latter approach avoids retraining but introduces key challenges: (i) Numerical encoding issues—LLM tokenizers, designed for text, struggle with numerical values, treating consecutive numbers as independent tokens (Nate Gruver and Wilson, 2023) and failing to preserve temporal dependencies (Spathis and Kawsar, 2024). (ii) Sequence length constraints—sensor data often exceeds LLMs’ maximum context length, leading to truncation, information loss, and increased computational costs. (iii) Multi-channel complexity—LLMs process univariate inputs, making it difficult to encode multi-sensor data in a way that retains inter-channel dependencies. (iv) Prompt engineering challenges—designing effective prompts that enable LLMs to interpret numerical sensor readings, detect trends and classify activities remains a challenge (Liu et al., 2023b).
To address these challenges, we propose SensorLLM, a framework that aligns sensor data with natural language, allowing LLMs to analyze sensor data through text-based interactions (see Figure 1) without modifying the LLM itself. A major challenge is the annotation bottleneck, further compounded by the complexity and heterogeneity of sensor data. Unlike image-text pairs, sensor data comprises multi-channel numerical signals with diverse characteristics, making direct interpretation and alignment difficult.
@@
Existing methods (Jin et al., 2024a; Sun et al., 2024a) have explored condensed text prototypes for alignment, but these approaches often lack interpretability and require extensive tuning to select suitable prototypes. In contrast, we propose an automatic text generation approach that aligns with human intuition by deriving descriptive trend-based text directly from time-series data using statistical analyses and predefined templates. This method is precise, scalable, and interpretable, eliminating the need for manual annotations while preserving essential sensor characteristics. SensorLLM follows a two-stage framework:
@@
We automatically generate question-answer pairs to align sensor data with text while preserving temporal features using a pretrained encoder. The resulting embeddings are mapped into a space interpretable by the LLM, mitigating issues associated with text-specific tokenization. Additionally, we introduce special tokens for sensor channels, enabling LLMs to effectively capture multi-channel dependencies.
@@
The aligned embeddings are utilized for HAR, leveraging the LLM’s reasoning capabilities while keeping its parameters frozen. This design extends LLMs beyond their original training, addressing concerns raised by Tan et al. (2024) regarding their applicability to time-series data. To our knowledge, this is the first approach to integrate sensor data into LLMs for sensor data analysis and HAR tasks.
To our knowledge, this is the first approach to integrate sensor data into LLMs for sensor data analysis and HAR tasks. The key contributions of this work are as follows:
  • We propose a fully automated, human-intuitive approach for aligning time-series data with descriptive text, eliminating the need for manual annotations. Using text similarity metrics, human evaluations, and LLM-based assessments, we demonstrate that SensorLLM effectively captures temporal patterns and channel-specific features, enabling robust multimodal understanding.
@@ 3 Methods
In this work, we propose SensorLLM, a framework that aligns sensor data with descriptive text through automatically generated question–answer pairs. Our goal is to develop a multimodal model with reasoning capabilities for analyzing wearable sensor data. As shown in Figure 2, SensorLLM comprises three core components: (1) a pretrained LLM, (2) a pretrained time-series (TS) embedder, and (3) an alignment module MLP, all operating in a two-stage framework.
In the Sensor-Language Alignment Stage, a generative model aligns sensor readings with text based on user instructions (Liu et al., 2023a) and questions. In the Task-Aware Tuning Stage, a classifier is added on top of the LLM to perform HAR. Crucially, only the MLP in both stages and the classifier in the Task-Aware Tuning Stage are trainable, while the backbone LLM and TS embedder remain frozen. This design yields a highly efficient and lightweight training process, requiring only 5.67% (535.9M) of the parameters to be trainable in the alignment stage and 0.12% (10.5M) in the tuning stage.
@@
3.1 Sensor-Text Data Generation
Aligning time-series data with text for LLM-based tasks is challenging due to the lack of rich semantic labels beyond class annotations, making manual annotation impractical (Deldari et al., 2024; Haresamudram et al., 2024). While prior works rely on predefined text prototypes (Sun et al., 2024b; Jin et al., 2024a), we aim for a more human-intuitive representation of sensor data.
We argue that time-series data inherently contains semantic patterns that can be expressed through descriptive text, from simple numerical trends to statistical insights. To achieve this, we automatically generate descriptive text by analyzing observed trends and fluctuations in the data. Using predefined templates, we construct diverse question-answer (QA) pairs that capture trend changes while ensuring accuracy and scalability. These templates (Appendix A.2) are randomly combined to enhance diversity. For example:
@@ 3.2 Sensor-Language Alignment
As shown in Figure 2 (a), the Sensor-Language Alignment stage employs a generative model to create multimodal sentences that combine single-channel sensor readings with textual descriptions. The sensor data is represented as a matrix $\mathbf{X}\in\mathbb{R}^{C\times T}$, where $C$ is the number of sensor channels and $T$ is the sequence length. Each channel’s data, denoted as $\mathbf{X}^{c}$ for channel $c$, is processed independently to preserve its unique characteristics. The data is then divided into non-overlapping segments, $\mathbf{X}_{S}^{c}$, where $S$ is the total number of segments. Each segment $x_{s}$ is assigned a random length $l$ within a predefined range, allowing the model to learn from varying temporal patterns and trend variations. This segmentation strategy ensures that both long-term trends and short-term fluctuations are effectively captured in the generated multimodal sentences.
We use Chronos (Ansari et al., 2024) as the TS embedder, generating segment embeddings $\hat{x}_{s}\in\mathbb{R}^{(l+1)\times d_{ts}}$, where $d_{ts}$ is the feature dimension, and $(l+1)$ accounts for the [EOS] token appended during Chronos tokenization. Prior to inputting sensor segments into Chronos (Appendix A.3), we apply instance normalization $\tilde{x}_{s}=\frac{x_{s}-\text{mean}(x_{s})}{\text{std}(x_{s})}$ to standardize the data. Llama3-8B (Touvron et al., 2023) serves as our LLM backbone.
@@
3.3 Task-Aware Tuning
Input Embedding.
For each sensor channel $c$, we retrieve its aligned sensor embeddings $\hat{o}_{s}^{c}$. These embeddings are then concatenated across all channels, along with their corresponding statistical features (mean and variance), to form the final input embedding:
where $\hat{z}_{\text{stat}}$ represents the statistical information, and $C$ is the number of channels. This ensures the model integrates both temporal and statistical characteristics for HAR.
Loss Function.
$\mathcal{L}_{cls}=-\sum_{i=0}^{M-1}y_{i}\log\hat{y}_{i},$ (4)
where $y_{i}$ is the ground truth label.
4 Experiments
In this section, we evaluate SensorLLM in enabling LLMs to interpret, reason about, and classify sensor data for HAR tasks. All experiments are conducted on NVIDIA A100-80G GPUs. To assess the LLM’s ability to learn and generalize from raw sensor inputs, we ensure that the same training and testing subjects are used in both the Sensor-Language Alignment and Task-Aware Tuning stages. This guarantees that test data in the second stage remains unseen during alignment, ensuring a fair evaluation of generalization. We select Chronos as the TS embedder because it has not been pre-trained on motion sensor data, making it an ideal candidate for evaluating our approach’s robustness in learning directly from unprocessed sensor signals.
4.1 Datasets
To evaluate the effectiveness and generalizability of SensorLLM, we conduct experiments on five publicly available HAR datasets: USC-HAD (Zhang and Sawchuk, 2012), UCI-HAR (Anguita et al., 2013), PAMAP2 (Reiss and Stricker, 2012), MHealth (Baños et al., 2014), and CAPTURE-24 (Chan et al., 2024). These datasets differ in sensor placement, sampling rates, channel configurations, and activity types, covering both controlled laboratory conditions and free-living environments. Additionally, they vary in scale, with USC-HAD, UCI-HAR, PAMAP2, and MHealth collected from a limited number of subjects, whereas CAPTURE-24 features a large-scale dataset with sensor recordings from 151 participants in real-world settings. Full dataset details, including subject count, sensor configurations, data splits, activity classes, preprocessing steps, and windowing strategies, are provided in Appendix A.6.
@@ 4.1 Datasets
Table 1: Evaluation of Sensor Data Trend Analysis Tasks for SensorLLM and GPT-4o. The assessment includes human and GPT-4o ratings (from 1 to 5, with 5 being the highest), as well as BLEU-1, ROUGE-1, ROUGE-L, METEOR, SBERT, and SimCSE (in %). The column GPT-4o refers to the trend analysis generated by GPT-4o itself, while the row GPT-4o refers to GPT-4o’s evaluation of the generated outputs.

### v3

@@ Abstract
We introduce SensorLLM, a two-stage framework that enables Large Language Models (LLMs) to perform human activity recognition (HAR) from wearable sensor data. While LLMs excel at reasoning and generalization, they struggle with time-series inputs due to limited semantic context, numerical complexity, and sequence variability. To address these challenges, we construct SensorQA, a question-answering dataset of human-intuitive sensor-text pairs spanning diverse HAR scenarios. It supervises the Sensor-Language Alignment stage, where the model aligns sensor inputs with trend descriptions. Special tokens are introduced to mark channel boundaries. This alignment enables LLMs to interpret numerical patterns, channel-specific signals, and variable-length inputs—without requiring human annotation. In the subsequent Task-Aware Tuning stage, we adapt the model for multivariate HAR classification, achieving performance that matches or exceeds state-of-the-art methods. Our results show that, guided by human-intuitive alignment, SensorLLM becomes an effective sensor learner, reasoner, and classifier—generalizing across varied HAR settings and paving the way for foundation model research in time-series analysis. Our codes are available at https://github.com/zechenli03/SensorLLM.
[Uncaptioned image] SensorLLM: Human-Intuitive Alignment of Multivariate Sensor Data with LLMs for Activity Recognition
@@ 1 Introduction
Human Activity Recognition (HAR) is a time-series (TS) classification task that maps sensor signals, such as accelerometer and gyroscope data, to human activities. Traditional models like LSTM (Guan and Plötz, 2017; Hammerla et al., 2016) and DeepConvLSTM (Ordóñez and Roggen, 2016) learn high-level features but are task-specific and struggle to generalize across different sensor configurations and activity sets. In contrast, Large Language Models (LLMs) (Han et al., 2021) have shown remarkable success in integrating diverse data types (Liu et al., 2023a; Wu et al., 2023b; Yin et al., 2023), including text and images.
@@
Enabling LLMs to process wearable sensor data (Jin et al., 2023) requires either (1) pretraining or fine-tuning on TS data (Zhou et al., 2023a), which demands substantial computational resources and is hindered by limited and imbalanced labeled data, or (2) leveraging zero-shot and few-shot prompting by converting sensor data into text (Kim et al., 2024; Ji et al., 2024). The latter approach avoids retraining but introduces key challenges: (i) Numerical encoding issues—language model tokenizers, designed for text, struggle with numerical values, treating consecutive numbers as independent tokens (Nate Gruver and Wilson, 2023) and failing to preserve temporal dependencies (Spathis and Kawsar, 2024). (ii) Sequence length constraints—complex numerical sequence often exceeds LLMs’ maximum context length, leading to truncation, information loss, and increased computational costs. (iii) Multivariate complexity—LLMs process univariate inputs, making it difficult to encode multivariate sensor data in a way that retains inter-channel dependencies. (iv) Prompt engineering challenges—designing effective prompts that enable LLMs to interpret numerical sensor readings, detect trends and classify activities remains a challenge (Liu et al., 2023b).
To address these challenges, we introduce SensorLLM, a human-intuition inspired framework that aligns numerical sensor data with natural language. Unlike image-text pairs, sensor data consists of multivariate signals with complex and varied patterns, making annotation particularly difficult. To tackle this, we release SensorQA, a question–answer (QA) dataset of aligned sensor-text pairs spanning diverse HAR tasks. SensorLLM is trained on this dataset to learn intuitive mappings between sensor readings and descriptive language, enabling LLMs to interpret sensor data through natural interactions—without any modification to the LLM architecture (see Figure 1).
@@
Existing methods (Jin et al., 2024a; Sun et al., 2024a) have explored condensed text prototypes for alignment, but these approaches often lack interpretability and require extensive tuning to select suitable prototypes. In contrast, we propose an automatic text generation approach that aligns with human intuition by deriving descriptive trend-based text directly from TS data using statistical analyses and predefined templates. This method is precise, scalable, and interpretable, eliminating the need for manual annotations while preserving essential sensor characteristics. SensorLLM follows a two-stage framework:
@@
Our SensorQA pairs each sensor window with QA pairs to align sensor features with natural language in a structured format. A pretrained TS encoder extracts temporal features from the sensor data, which are then projected into a space interpretable by the LLM, mitigating text-specific tokenization issues. We also introduce special tokens for each sensor channel, enabling LLMs to effectively capture multivariate (multi-channel) dependencies.
@@
The aligned embeddings are used for HAR, leveraging the LLM’s reasoning capabilities while keeping its parameters frozen. Importantly, our framework naturally supports sensor inputs of varying sequence lengths and arbitrary numbers of channels (i.e., multivariate TS data), a flexibility that prior approaches have struggled to accommodate.
To the best of our knowledge, this is the first approach to integrate sensor data into LLMs for sensor-based analysis and activity recognition. The key contributions of this work are:
  • We release SensorQA, a novel human-intuitive dataset of aligned sensor-text QA pairs covering diverse HAR scenarios. We evaluate SensorLLM through text similarity metrics, human judgments, and LLM-based assessments, confirming its ability to capture temporal patterns for robust multimodal understanding. SensorQA and SensorLLM support sensor inputs with varying sequence lengths and channel configurations, allowing broad and realistic HAR evaluation.
@@ 3 Methods
In this work, we propose SensorLLM, a two-stage framework that aligns wearable sensor data with descriptive text by a high-precision corpus of question-answering pairs created without any human annotations and tailored for wearable sensor reasoning. Our aim is to build a multimodal model capable of interpreting and reasoning over time-series (TS) signals. As shown in Figure 2, SensorLLM consists of three core components: (1) a pretrained LLM, (2) a pretrained TS embedder, and (3) a lightweight MLP alignment module.
In the Sensor–Language Alignment stage, a generative model aligns sensor readings with text, and in the Task–Aware Tuning stage, a lightweight classifier is added on top of the LLM to perform HAR. Crucially, only the alignment MLP and this classifier are trainable—both the backbone LLM and the TS embedder remain frozen—resulting in just 5.67% (535.9 M) of parameters being fine-tuned in the first stage and 0.12% (10.5 M) in the second, making training extremely efficient.
@@
3.1 SensorQA Dataset
Aligning TS data with natural language is challenging due to the lack of rich semantic annotations beyond class labels, making manual labeling costly and impractical (Deldari et al., 2024; Haresamudram et al., 2024). While prior works often rely on fixed text prototypes (Sun et al., 2024b; Jin et al., 2024a), we introduce SensorQA —a scalable and human-intuitive dataset designed to bridge sensor data and language through structured QA pairs.
SensorQA is built on the idea that TS signals naturally exhibit semantic patterns—such as trends and statistical behaviors—that can be described in natural language. Using predefined templates (Appendix A.2), descriptive QA pairs are automatically generated for each sensor window without any human annotations. Each pair includes information such as sensor type, time range, and observed trends. Templates are randomly combined to enhance diversity. For example:
@@ 3.2 Sensor-Language Alignment
As shown in Figure 2 (a), the Sensor–Language Alignment stage uses a generative model to create multimodal sentences by combining single-channel sensor readings with text from our SensorQA dataset. The sensor data is represented as a matrix $\mathbf{X}\in\mathbb{R}^{C\times T}$, where $C$ is the number of channels and $T$ is the sequence length. Each channel’s data, denoted as $\mathbf{X}^{c}$, is processed independently to retain channel-specific characteristics. The data is segmented into non-overlapping segments $\mathbf{X}_{S}^{c}$, where $S$ is the total number of segments. Each segment $x_{s}$ is assigned a random length $l$ within a predefined range, encouraging the model to learn from both short-term fluctuations and long-term trends.
We use Chronos (Ansari et al., 2024) as the TS encoder to generate segment embeddings $\hat{x}_{s}\in\mathbb{R}^{(l+1)\times d{ts}}$, where $d_{ts}$ is the embedding dimension and $(l+1)$ accounts for the [EOS] token added during Chronos tokenization (Appendix A.3). Before feeding segments into Chronos, we apply instance normalization: $\tilde{x}_{s}=\frac{x_{s}-\text{mean}(x_{s})}{\text{std}(x_{s})}$. For the language backbone, we use LLaMA3-8B (Touvron et al., 2023).
@@
Table 1: Evaluation of Sensor Data Understanding tasks. The column GPT-4o denotes trend descriptions generated by GPT-4o, while the row GPT-4o indicates evaluations conducted by GPT-4o on the model outputs.
3.3 Task-Aware Tuning
Input Embedding.
For each sensor channel $c$, we retrieve its aligned embeddings $\hat{o}_{s}^{c}$ from the pretrained alignment module. These are concatenated across all channels, along with statistical features (mean and variance) from our SensorQA dataset, to form the final input embedding:
where $\hat{z}_{\text{stat}}$ represents the statistical information. This ensures the model integrates both temporal and statistical characteristics for HAR.
Loss Function.
$\mathcal{L}_{cls}=-\sum_{i=0}^{M-1}y_{i}\log\hat{y}_{i},$ (4)
where $y_{i}$ is the ground truth label.
Method USC-HAD UCI-HAR PAMAP2 MHealth CAPTURE-24
F1-macro Accuracy F1-macro Accuracy F1-macro Accuracy F1-macro Accuracy F1-macro Accuracy
PatchTST 45.2±$\pm 1.48$ 45.6±$\pm 2.19$ 86.8±$\pm 0.84$ 86.0±$\pm 0.71$ 82.0±$\pm 0.71$ 81.2±$\pm 0.84$ 80.0±$\pm 1.58$ 79.4±$\pm 1.34$ 35.6±$\pm 0.89$ 66.2±$\pm 1.10$
Ns-Transformer 52.6±$\pm 2.30$ 51.8±$\pm 2.86$ 88.0±$\pm 0.71$ 87.4±$\pm 0.55$ 78.8±$\pm 0.84$ 78.8±$\pm 0.84$ 77.2±$\pm 1.48$ 75.8±$\pm 1.48$ 34.8±$\pm 1.10$ 65.4±$\pm 0.55$
Informer 51.2±$\pm 1.30$ 51.6±$\pm 1.52$ 86.6±$\pm 1.14$ 86.4±$\pm 0.89$ 78.0±$\pm 1.58$ 78.6±$\pm 1.34$ 74.0±$\pm 0.71$ 72.8±$\pm 0.84$ 35.6±$\pm 0.55$ 66.8±$\pm 0.84$
Transformer 49.6±$\pm 1.67$ 50.6±$\pm 0.55$ 85.4±$\pm 0.89$ 85.2±$\pm 1.10$ 77.0±$\pm 0.71$ 77.6±$\pm 0.89$ 75.2±$\pm 1.30$ 74.6±$\pm 1.34$ 32.8±$\pm 0.84$ 65.4±$\pm 0.89$
iTransformer 48.4±$\pm 1.82$ 49.6±$\pm 1.67$ 81.8±$\pm 0.84$ 81.8±$\pm 0.84$ 76.6±$\pm 0.55$ 75.8±$\pm 0.45$ 80.4±$\pm 1.14$ 80.0±$\pm 1.22$ 19.8±$\pm 0.84$ 62.4±$\pm 0.89$
TimesNet 52.2±$\pm 2.39$ 52.6±$\pm 2.07$ 87.4±$\pm 1.14$ 86.6±$\pm 1.14$ 76.2±$\pm 1.92$ 77.4±$\pm 1.14$ 78.4±$\pm 1.52$ 77.2±$\pm 1.48$ 34.8±$\pm 0.84$ 65.8±$\pm 1.79$
GPT4TS 54.2±$\pm 2.05$ 56.0±$\pm 1.58$ 88.2±$\pm 0.84$ 87.6±$\pm 0.55$ 80.4±$\pm 0.89$ 79.8±$\pm 0.45$ 76.4±$\pm 1.14$ 75.4±$\pm 1.14$ 32.8±$\pm 1.10$ 62.2±$\pm 1.92$
Chronos+MLP 44.2±$\pm 1.30$ 44.0±$\pm 0.71$ 82.2±$\pm 0.84$ 81.2±$\pm 0.84$ 79.8±$\pm 0.45$ 79.8±$\pm 0.45$ 83.0±$\pm 0.71$ 82.0±$\pm 0.71$ 38.0±$\pm 0.71$ 68.2±$\pm 0.84$
DeepConvLSTM 48.8±$\pm 2.39$ 50.6±$\pm 2.41$ 89.2±$\pm 0.84$ 89.2±$\pm 0.84$ 78.4±$\pm 1.52$ 78.2±$\pm 1.10$ 75.0±$\pm 1.87$ 76.0±$\pm 1.00$ 40.4±$\pm 0.89$ 69.4±$\pm 1.14$
DeepConvLSTMAtt 54.0±$\pm 2.12$ 54.4±$\pm 3.21$ 89.6±$\pm 1.14$ 89.4±$\pm 1.14$ 79.2±$\pm 1.30$ 79.6±$\pm 1.14$ 77.4±$\pm 2.19$ 76.8±$\pm 1.48$ 41.4±$\pm 0.55$ 70.4±$\pm 0.55$
Attend 60.2±$\pm 2.17$ 60.8±$\pm 1.92$ 93.2±$\pm 0.84$ 92.8±$\pm 0.45$ 84.6±$\pm 1.14$ 85.0±$\pm 0.71$ 83.4±$\pm 1.14$ 82.6±$\pm 1.14$ 43.6±$\pm 0.55$ $\pm 0.71$
SensorLLM 61.2±$\pm 3.56$ 62.6±$\pm 3.36$ 91.2±$\pm 1.48$ 90.8±$\pm 1.30$ 86.2±$\pm 1.48$ 87.2±$\pm 0.84$ 89.4±$\pm 3.85$ 89.0±$\pm 3.54$ 48.6±$\pm 1.14$ 72.0±$\pm 0.71$
Table 2: F1-macro and accuracy scores (%) for the Human Activity Recognition tasks, presented as the mean and standard deviation over 5 random repetitions. Bold for the best and underline for the second-best.
4 Experiments
In this section, we evaluate SensorLLM in enabling LLMs to interpret, reason about, and classify sensor data for HAR tasks. All experiments are conducted on NVIDIA A100-80G GPUs. To assess the LLM’s ability to learn and generalize from raw sensor inputs, we ensure that the same training and testing subjects are used in both the Sensor-Language Alignment and Task-Aware Tuning stages. This guarantees that test data in the second stage remains unseen during alignment, ensuring a fair evaluation of generalization. We select Chronos as the TS embedder since it has not been pre-trained on HAR-specific data, allowing us to evaluate the robustness of our approach in adapting to raw, domain-agnostic sensor signals.
4.1 Datasets
To evaluate the effectiveness and generalizability of SensorLLM, we conduct experiments on five publicly available HAR datasets: USC-HAD (Zhang and Sawchuk, 2012), UCI-HAR (Anguita et al., 2013), PAMAP2 (Reiss and Stricker, 2012), MHealth (Baños et al., 2014), and CAPTURE-24 (Chan et al., 2024). These datasets vary widely in subject counts, sensor placement, sampling rates, channel configurations, and activity types, covering both controlled laboratory conditions and free-living environments. SensorQA (see Appendix A.7) is built on these five benchmarks and supports both alignment and activity classification. All datasets are publicly available, containing no personally identifiable information, thus posing minimal ethical or privacy concerns.
We use subject-independent splits for all datasets except UCI-HAR, which comes with a fixed split. In all other datasets, training and test sets come from different subjects, ensuring the model is evaluated on unseen users. Full dataset details, including subject count, sensor configurations, data splits, activity classes, preprocessing steps, and windowing strategies, are provided in Appendix A.6.